OpenStudy (anonymous):

Hi all! This one has been troubling me. Solve the inequality: av(x-1)-av(x-3) >or equal to 5

ganeshie8 (ganeshie8):

\(\large a^{x-1} - a^{x-3} \ge 5\)

ganeshie8 (ganeshie8):

is it like above ?

OpenStudy (anonymous):

umm let me try to make the equation

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

\[\left| x-1 \right|-\left| x-3 \right| \ge 5\]

OpenStudy (anonymous):

it just keeps cancelling out when I do it, so I want to be sure I'm not doing something wrong

ganeshie8 (ganeshie8):

okay ! its a combination of two absolute value functions

ganeshie8 (ganeshie8):

at wat point u got stuck ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so I know it can be positive/negative, but when when I solve I get an answer that doesnt make sense like 2 is greater than/equal to 5

ganeshie8 (ganeshie8):

good :) that means, there are no solutions for this inequality

ganeshie8 (ganeshie8):

wat u arrived at is correct, there are no solutions for this guy

OpenStudy (anonymous):

aha, I just get paranoid when there isn't a solution XD

ganeshie8 (ganeshie8):

lol who doesnt ! btw how did u solve it ?

OpenStudy (anonymous):

umm well I had two solutions

ganeshie8 (ganeshie8):

did u square both sides ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I erased it let me rewrite real quick

ganeshie8 (ganeshie8):

its okay :) just wondering wat method u used. i did this by squaring both sides, to get rid of absolute bars

OpenStudy (anonymous):

If x>0 2>or equal to 5 for example

OpenStudy (anonymous):

then for x<0

OpenStudy (anonymous):

what is the squaring method?

ganeshie8 (ganeshie8):

\(\left| x-1 \right|-\left| x-3 \right| \ge 5 \) \(\left| x-1 \right|~~~\ge~~~ 5 +\left| x-3 \right| \)

ganeshie8 (ganeshie8):

square both sides

ganeshie8 (ganeshie8):

\(\left| x-1 \right|-\left| x-3 \right| \ge 5 \) \(\left| x-1 \right|~~~\ge~~~ 5 +\left| x-3 \right| \) \((x-1)^2~~~\ge~~~ 25 + (x-3)^2 + 10\left| x-3 \right| \)

OpenStudy (anonymous):

I see interesting

ganeshie8 (ganeshie8):

if u simplify and keep going, u wil arrive at 2=3 or coconuts = oranges kindof statement so we conclude NO solution exist

OpenStudy (anonymous):

cool! I didn't know you could do that

OpenStudy (anonymous):

Thanks again for you help as usual :))

ganeshie8 (ganeshie8):

\(\left| x-1 \right|-\left| x-3 \right| \ge 5 \) \(\left| x-1 \right|~~~\ge~~~ 5 +\left| x-3 \right| \) \((x-1)^2~~~\ge~~~ 25 + (x-3)^2 + 10\left| x-3 \right| \) .... \( 4x-33 ~~~\ge~~~ 10\left| x-3 \right| \) square both sides again ....

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

^_^